# BRiTE: Bootstrapping Reinforced Thinking Process to Enhance Language Model Reasoning

### v1

@@ B.2 Additional Theoretical Results
Proof of Theorem B.3.
By the same derivation of (B.1), we have
@@
$\qquad-\left\langle\nabla A(\theta_{t+1})-\nabla A(\theta_{t}),f_{\theta_{t}}-f_{\theta_{t+1}}\right\rangle,$ (B.19)
where the last equality is implied by the optimality condition in (B.8). By Lemma B.1, we have
$\mathrm{KL}\big{(}\mathbb{P}(\cdot,\cdot{\,|\,}x,\theta_{t})\|\mathbb{P}(\cdot,\cdot{\,|\,}x,\theta_{t+1})\big{)}=A(x,\theta_{t+1})-A(x,\theta_{t})+\langle\nabla A(x,\theta_{t}),f_{\theta_{t}}-f_{\theta_{t+1}}\rangle.$ (B.20)
Combining (B.2) and (B.20), we have
@@
Together with Lemma B.1, we have
@@
which finishes the proof of Theorem B.3. ∎
Appendix C Implementation Details
C.1 Implementation Details for Math Task
@@
C.2 Implementation Details for Coding Generation Tasks
@@
For the code generation task, We choose the first 4000 rows of the educational instruct split of the dataset OpenCoder-LLM/opc-sft-stage2444https://huggingface.co/datasets/OpenCoder-LLM/opc-sft-stage2/tree/main(Huang et al., 2024) as the training dataset, which comprises (instruction, code, test case) tuples. The entire dataset is generated with an algorithmic corpus as a seed and validated through a Python compiler.

### v2

@@ C.2 Additional Theoretical Results
Proof of Theorem C.3.
By the same derivation of (C.1), we have
@@
$\qquad-\left\langle\nabla A(\theta_{t+1})-\nabla A(\theta_{t}),f_{\theta_{t}}-f_{\theta_{t+1}}\right\rangle,$ (C.19)
where the last equality is implied by the optimality condition in (C.8). By Lemma C.1, we have
$\mathrm{KL}\big{(}\mathbb{P}(\cdot,\cdot{\,|\,}x,\theta_{t})\|\mathbb{P}(\cdot,\cdot{\,|\,}x,\theta_{t+1})\big{)}=A(x,\theta_{t+1})-A(x,\theta_{t})+\langle\nabla A(x,\theta_{t}),f_{\theta_{t}}-f_{\theta_{t+1}}\rangle.$ (C.20)
Combining (C.2) and (C.20), we have
@@
Together with Lemma C.1, we have
@@
which finishes the proof of Theorem C.3. ∎
Appendix D Implementation Details
D.1 Implementation Details for Math Task
@@
D.2 Implementation Details for Coding Generation Tasks
@@
For the code generation task, We choose the first 4000 rows of the educational instruct split of the dataset OpenCoder-LLM/opc-sft-stage2555https://huggingface.co/datasets/OpenCoder-LLM/opc-sft-stage2/tree/main(Huang et al., 2024) as the training dataset, which comprises (instruction, code, test case) tuples. The entire dataset is generated with an algorithmic corpus as a seed and validated through a Python compiler.
@@
D.3 Implentation Details for the scaling-up Experiments
Implementation of BRiTE.
We train the BRiTE algorithm on eight NVIDIA H100 GPUs using the GRPO pipeline (Shao et al., 2024) in the verl codebase (Sheng et al., 2024), learning the sampling policy $\mathbb{Q}$ (3.4) with a learning rate of $61e-$ 6 and a batch size of $512\times 8$ (8 rollouts per prompt). For the subsequent SFT $\psi$-update on rationales sampled by $\mathbb{Q}$, we set the learning rate to $61e-$ 6 and bath size to be $8$. We also adapt LoRA and set r to be $32$ and lora alpha to be $128$.
Implementation of RS.
For rejection sampling, we set the temperature of the model’s generation to $1.0$, sample $N=2$ candidate rationales for each problem $x$, and select the best rationales $z_{\text{rs}}$. we filter generations by selecting those that produce the correct final answer. The model is then fine-tuned on these problem-rationale-answer tuples $\{x,z_{\text{rs}},y^{*}\}$ using the same learning rate and LoRA parameters as in BRiTE.